# Annotating Ambiguous Images: General Annotation Strategy for Image Classification with Real-World Biomedical Validation on Vertebral Fracture Diagnosis

### v1

@@ 3.2 Analysis
Table 2: Comparison human and network performance on Verse2019 [35] – * based on ten random sampled annotations
A comparison of the annotator and network performance on the Verse2019 [35] test set is given in Table 2. When comparing the macro F1 scores in 2D between annotations performed with and without proposals and those with proposal-based guidance achieved a significantly higher F1 score of about 2%. This means that proposal-based guidance led to better annotations by the annotators. The average macro F1 score for 3D annotations shows no significant difference compared to the 2D annotations, meaning that the modality difference was of no to little importance to the annotators.
@@
The best performing network on the validation data for the proposal generation did not surpass the set target threshold $\mu=0.6$ with a macro F1 of 0.57. However, it surpassed $\mu$ on the validation and training data and it was decided that the performance was close enough. The proposal network was a semi-supervised Mean-Teacher model [53] with overclustering and default hyperparameters from [46]. The performance of the model was significantly better when trained only on the mask (see Figure 3) with only the desired vertebra visible instead of all other 2D modalities. We attribute this difference to the fact that the model could be more easily confused with vertebrae that are not currently of interest and to the more complicated representation of information e.g. in the form of blended images. In terms of the performance of the network trained on our new annotated data, but evaluated on the original test ground truth, we found that more annotations sampled from the annotated data generally led to better performance. This sampling from the human annotation could achieve random peaks up to about 0.7 F1-score. A standard ResNet50 model [18] with default hyperparameters (see supplementary for details) and cross-entropy as loss function could achieve a score of 0.63, which is higher than the more complex proposal network (trained on the original training data) and previously reported values by Wei et al. for non-specialized networks [59]. While they show that they can achieve higher scores with more problem-specific loss functions, with our data-centric approach, we can achieve an improvement without changing the model or training purely from the input data. Future potential improvement to also use special loss functions and better backend-models non-withstanding.
method number of annotations
proposal blending correction 1 3 5 10
Table 3: Resulting KL across different methods and number of annotations – The first method is the recommendation based on our defined guidelines. The next two methods use proposals and the last three methods do not use proposals which means their annotation time is also slower in comparison to the recommendation. Thus, one should not compare with the same number of annotations but higher number of annotations for DC3 proposals because the higher number can be achieved potentially in the same or less time. The results are given as KL $\pm$ STD ( Relative Change in % with regard to recommended method)
In Table 3, we compare the network predictions with the approximated soft ground truth distribution ($P(L^{x}=\cdot)$) using the Kullback-Leibler-divergence [27]. Comparing the use of DC3 proposals with and without blending or correction, we see that the results are better with both improvement methods proposed in [44]. Thus, we can confirm that the application of CleverLabel (the combination of both improvements) should be done in the post-processing step. We can also confirm that applying balanced blending to annotations without proposals improves the results. Using only the class blending based on the majority class seems to work particularly well for this task. However, as the number of annotations increases, the benefit decreases and from 5 annotations onward, it becomes worse than the recommended method of our defined strategy. In our description of how to apply the proposed annotation strategy (see subsection 3.1), we stated that the introduction of a bias is not a major concern. Nevertheless, we can validate that the strategy leads to the best results even if a bias would not be acceptable. Based on the speedup of about 1.2, the recommendation would have been to not annotate with proposals and to use only balanced blending, which is the best method for multiple annotations. Theoretically, if we had a high speedup, we could compare a higher number of annotations with proposals to fewer annotations without proposals. For example, a speedup at the decision boundary of 3, 5 annotations with proposal and 3 annotations without proposal, we have a lower annotation cost based on the calculations form [44] (5 $\cdot$ 0.33 vs. 3) and have a lower KL score. Thus, even in the case of higher speedups, our strategy suggests the correct choice.
4 Conclusion
In conclusion, this study presents a general strategy for annotating and processing ambiguous images with verification in vertebral fracture diagnosis. Overall, the human performance analysis demonstrated the effectiveness of proposal-guided annotation, resulting in improved F1 scores and reduced training time. A neural network trained on newly annotated data achieved a higher classification performance on the original test set and can potentially more than halve the KL score on our new baseline. Based on the results for all possible scenarios, we were able to verify the effectiveness of our strategy based on the literature for the given use case. By providing practical guidance and verifying its application on a real-world vertebral fracture dataset, this research contributes to the creation of high-quality datasets and advances the field of data-centric deep learning.
Supplementary Material
Approximation of workload
We look at the expected number of annotations per hour done by one annotator as $\frac{a}{h}$. Depending on the expected portion of complete consensus of the dataset $p_{c}$ and the total dataset size $N=|X_{u}|$, we can calculate the expected annotation time in hours with $A_{N}=N\cdot(p_{c}\cdot A_{c}+(1-p_{c})\cdot A^{c})/\frac{a}{h}$. In Figure 4 the required annotation time $A_{N}$ in days for $N=10,000,A_{c}=10,A^{c}=50$ with varying $p_{c}$ is given as well as values reported in [46]. The reported values for $\frac{a}{h}$ are between 1,000 and 4,500 without proposals or 3,000 and 10,000 with proposals. Thus, even in the worst-case the expected workdays with proposals are less than 10 days, or two weeks, which should be feasible in most cases especially when multiple annotators are available. These values are examples and have of course to be adopted to the specific use-case.
Figure 4: Annotation time in days based on expected consensus ratio $p_{c}$ and annotations per hour $\frac{a}{h}$, reported values from [46] and with 95% agreeing votes threshold $\hat{p_{c}}$
Approximation of confidences
Considering the binary case of the annotation task for class $k$ for each image $x$, i.e. whether class $k$ is annotated for $x$ or not, and having a sufficient number of annotations $A$, the 95%-confidence interval can be calculated using the Wald confidence interval [37]. The interval is given by
$\begin{split}P(L^{x}=k)&\pm Z_{95\%}\cdot\sqrt{\frac{p_{k}(1-p_{k})}{A}}\end{split}$ (1)
with 95%-confidence value, which is 1.96, and $p_{k}$ is the unknown true GT probability for class $k$. The confidence interval is the largest for $p=0.5$. The formula can be rearranged to calculate the required annotations $A$ if a 95% confidence interval around the approximation $P(L^{x}=k)$ of width $W$ and margin $W/2$ is to be guaranteed.
Figure 5 visualizes the required annotations $A$ for various $p$ and $W$. These calculations can guarantee with 95% confidence that the confidence interval has a width of $W\approx 1$ for $A=3$, $W\approx 0.62$ for $A=10$ and $W\approx 0.28$ for $A=50$. However, a high number of $A$ is indirectly assumed because the Wald confidence interval approximates the original binomial distribution with a normal distribution. This approximation is especially inaccurate for small numbers of $A$ and $p$ near 0 or 1, as pointed out by Lyman et al. [37]. They suggest to approximate the 95% confidence intervals with $((0.25)^{\frac{1}{A}},1)$ when $p$ is close to one. This would give a lower bound of the interval of $0.63$ for $A=3$, $0.87$ for $A=10$ and $0.97$ for $A=50$. All these calculations assume an idealized version of the annotators, which must be taken into account in their interpretation. For this reason, it will not be investigated further how the confidence intervals can be approximated more precisely, but conclude that a consensus decision is uncertain for $A=3$ annotations, but reasonable for ten or more annotations. Similarly, approximations of $p$ with a confidence interval width less than 0.3 are only reasonable for about 50 annotations.
Figure 5: Number of images for confidence width $W$, logarithmic scale
Additional minor considerations for Annotation Strategy
  • Constraints like collaboration modes and support for proposals should be considered. For example, if crowd-sourcing should be used for annotation, this restricts the appropriate platforms. Due to the fact if proposals will be used is decided later in subsection 2.3, it is advisable to ensure that these options are supported to avoid unnecessary backtracking.
  • Proposals $\rho_{x}$ with a low ground truth probability ($P(L^{x}=\rho_{x})=0$) should be avoided and [46] indicates only up to 10% of all proposals should be of such poor quality to allow successful training. The distribution $P(L^{x}=\cdot)$ might not be known at this point but if available this should be checked in addition to the threshold $\mu$.
  • The variables $c,\delta$ and $S$ may be approximated on various ways. One possible approach is defined in [44] by annotating about 100 images at least 10 times for the estimation of $c$, $\delta$ and $P(L^{x}=\cdot)$. The speedup $S$ can for example be measured on a small subset with and without proposals during the training of the annotators. However, if the variables are not required during the decision process described above, they could also be approximated on the real annotation data or ignored entirely. For example, $c$, can easily be approximated on the complete dataset if no proposals are used during the annotation process and an approximation of $S$ is not required if a small introduction of a bias is acceptable. The variables could also be approximated with an educated guess either based on previous knowledge or experiments.
  • The question is how $A_{c},A^{c}$ should be selected. We concentrate on two major aspects, what is a feasible workload for the annotators and what guarantees can be given. See for more details the approximations in Approximation of workload and Approximation of confidences.
  • The blending step may not be necessary if the number of annotations is high (e.g. above 50), since the distribution is already approximated quite accurately.
Approximation of $\delta$
We used annotations created during the training of the annotators to estimate $\delta$. We randomly picked 20 examples without consensus from the annotations without proposals. Based on these samples, the proposal per image $p_{x}$, the probabilities $P(L_{x}=\rho_{x})$ with and without proposal and the formula for simulated proposal acceptance of [44], we calculated $\delta\approx 0.11$.
Used hyperparameters
The used hyperparameters are defined in Table 4. For more detail see the respective modified source code.
Parameter Proposal network Evaluation network
Architecture DC3[46] + Mean-Teacher[53] ResNet50[18]
Learning rate 0.03 0.1
Epochs 600$\dagger$ 60
Optimizer SGD SGD
Batch size 64 128
Weight Decay 0.0005 0.0005
Pretrained Weights N/A ImageNet
Source code [46] [45]
Additional results of the learning effect
Regarding the learning effect of the annotators, the training iterations showed an increase in F1 scores over time. The improvement in F1 scores across all annotators between the training and the last three iterations was 6.79 without proposals and 9.19 with proposals. When comparing the benefit of DC3 proposals versus no proposals, the average improvement in F1 scores was 3.27 for the first three iterations and 0.41 for the last three iterations. The proposal still has in the end a benefit but the annotators can more successfully annotate event without them. Moreover, the introduction of proposal-based annotations resulted in a reduction in training time which also diminished over time. The first three iterations saw a reduction of 13.49 minutes, while the last three iterations experienced a decrease of 4.55 minutes. Overall, we can conclude that the proposals helped to improve the quality and reduce the annotation time. However, this effect was diminished overtime while the performance with and without proposals increase. Thus, we see a learning effect of the annotators.
Annotation platform
Figure 6 displays the used annotation website. It is important to note, that all images are displayed to the user in a grid. In the case without proposal (see Figure 6), the shown classes are mixed and all have to be annotated manually by the user. In the second case with proposal (see Figure 6), most of the images are of the same class and can be annotated simultaneously, while only the errors need to be corrected.
[…]
Figure 6: Screenshots of the annotation website on the MiceBone dataset from [45], The sidebar for data selection is hidden for data protection reasons and to keep the anonymity.

### v2

@@ E.2 Evaluation of human performance
A comparison of the annotator and network performance on the Verse2019 (Löffler et al., 2020) test set is given in Table 3. When comparing the macro F1 scores in 2D between annotations performed with and without proposals and those with proposal-based guidance, the latter achieved a significantly higher F1 score of about 2%. This means that the employment of proposal-based guidance resulted in improved annotations by the annotators. The average macro F1 score for 3D annotations shows no significant difference compared to the 2D annotations, meaning that the modality difference was of no to little importance to the annotators as expected in subsection 3.1.
@@
Appendix F Dataset Card
Attribute Details
Dataset Name VerSe 2D Projections with Soft Labels
Dataset URL https://zenodo.org/records/8115942
Original datasets:
VerSe’19: https://osf.io/nqjyw/ VerSe’20: https://osf.io/t98fz/
Description The dataset contains 2D projection of the central slices for vertebra of the VerSe (Large Scale Vertebrae Segmentation Challenge) datasets. These are collections of multi-detector CT scans for vertebral labelling and segmentation, significant for its size, diversity, and detailed annotations. The data, collected in conjunction with MICCAI 2019 and 2020, is anonymized and includes a variety of normal and pathological cases.
Data Collection Domain: Medical Imaging, 2D Projections of Spine Multi-detector Computed Tomography (MDCT)
Annotation Details: Soft-labels by evaluating proposal-based human non-expert annotations
VerSe’19: 160 CT scans from 141 patients, 1725 annotated vertebrae
VerSe’20: 374 CT scans from 335 patients, 4505 annotations
Data Preprocessing Scans and segementation masks have been anonymized and converted to NIFTI format (https://nifti.nimh.nih.gov/) in the VerSe dataset. The resolution of the CT images was adjusted to limit computational demands for deep learning algorithms. Coordinates of vertebral body centroids per vertebral level are stored in JSON format.
Annotations The proposed multi-step algorithmic framework has been applied to generate annotations with or without proposals by trained non-experts, including post-processing with CleverLabel and class blending.
Usage Development and evaluation of algorithms for vertebral labelling and segmentation in spine CT scans. VerSe‘19 and VerSe‘20 were used in challenges to evaluate accuracy, precision, and the effectiveness of segmentation and fracture detection as part of MICCAI 2019 and 2020.
Known Limitations While including multiple sites, the datasets might still not represent all possible variations in spine pathology and patient demographics.
VerSe’19: Exclusion of patients younger than 30 years and certain spinal conditions like bone metastasis
Ethical Considerations Both VerSe studies were approved by the local institutional review board, which conducted a thorough review of the research methodology, ensuring adherence to ethical guidelines. All patient data included in this study were anonymized to protect the privacy and confidentiality of the individuals.
Acknowledgements VerSe has been supported by the European Research Council under Horizon2020 (GA637164–iBack–ERC–2014–STG) and NVIDIA.
Table 4: Overview of the VerSe 2D Projections with Soft Labels